# Density Estimation via Measure Transport: Outlook for Applications in the Biological Sciences

### v2

@@ Abstract
One among several advantages of measure transport methods is that they allow for a unified framework for processing and analysis of data distributed according to a wide class of probability measures. Within this context, we present results from computational studies aimed at assessing the potential of measure transport techniques, specifically, the use of triangular transport maps, as part of a workflow intended to support research in the biological sciences. Scarce data scenarios, which are common in domains such as radiation biology, are of particular interest. We find that when data is scarce, sparse transport maps are advantageous. In particular, statistics gathered from computing series of (sparse) adaptive transport maps, trained on a series of randomly chosen subsets of the set of available data samples, leads to uncovering information hidden in the data. As a result, in the radiation biology application considered here, this approach provides a tool for generating hypotheses about gene relationships and their dynamics under radiation exposure.
@@ 1 Introduction
The problem of estimating a probability distribution density from samples (e.g., observations, measurements, or simulation data) is ubiquitous in data science, uncertainty quantification, clustering and classification, and probabilistic modeling and inference tasks. Moreover, it is common among various scientific and engineering fields, including biology [38, 14, 1, 41, 5, 7, 12, 39]. Often, well-known parametric density functions (dependent on few parameters), such as the Gaussian or Weibull density distribution functions, are adopted. While this may simplify certain tasks (e.g., computational ones), many of these known density distribution functions are not necessarily suitable for characterizing data that exhibit complex features, such as (spatial and/or temporal) correlations and non-Gaussian characteristics. For instance, as reported in [7], accounting for differences in the distribution densities of gene expressions can lead to improved interpretation of cancer transcriptomic data. Hence, a density estimation framework capable of characterizing a diverse range of properties is highly desirable. A measure transport approach [44, 37, 36] offers this possibility.
@@
Optimal measure transport, broadly defined, deals with the problem of minimizing the cost of transporting one (probability) measure to another. The subject has a long history with significant impact in many fields of mathematics, such as probability theory, optimization, and partial differential equations, among others [44, 37]. Current interest on the topic stems from its potential in numerous areas, including density estimation, data science, machine learning, Bayesian inference, sampling, and data assimilation [1, 42, 41, 24, 8, 40, 36, 32, 11, 39, 45, 17]. Among various measure transport techniques under active research, we mention the use of triangular transport maps, normalizing/gradient flows, diffusion maps, and invertible neural networks. Although differing in the particulars, affording a unified framework for processing and analysis of data from a range of distributions (Gaussian, non-Gaussian, uni-modal, multi-modal, etc.) is a powerful property common to all.
In this paper we consider measure transport for density estimation from data, in particular the use of triangular transport maps, envisioning its potential within a computational workflow for tasks associated with biological applications. The motivating problem, which is density estimation (or prior/posterior construction) when data is scarce, as well as the proposed workflow, are described in Section 2. Overviews of the topics of measure transport and density estimation via triangular transport maps appear, respectively, in Sections 3 and 4. References for thorough treatments of these subjects are provided therein. Computational experiments performed using gene expression data are presented in Section 5.
@@ 2 Motivating Application
The present work stems from the conception of a comprehensive computational workflow intended to support research in the biological sciences, in particular the radiation biology domain. Within this effort, our interest is in enabling a unified framework where probabilistic modeling, inference, and statistical analyses can be carried out for a wide range of data distributions. As noted in the Introduction (Section 1), measure transport offers the possibility of such unified approach. We envision measure transport techniques comprising one component within the larger scope, as depicted in Figure 1.
@@
Figure 1: Computational workflow for probabilistic modeling, inference, and statistical analyses. To account for scarcity of data, model training can be augmented with prior knowledge. Such prior knowledge may come from large language models or human experts, for instance. Measure transport techniques allow for a unified framework for dealing with data exhibiting diverse characteristics.
@@
One motivating problem within the framework from Figure 1 is that of probability distribution density estimation for phenotypic classification problems, for example, for inferring class labels, such as radiation exposure levels or cancer types, given sample data. Normalizing flows (one kind among the various measure transport methods) have been successfully used for clustering and classification in the cancer domain [1]. Now, when data is scarce, regardless of the density estimation technique adopted, prior knowledge (associated with the particular application) may be used in conjunction with data to estimate the class-conditional densities. In the work [5], the authors proposed incorporation of prior knowledge from biological pathways for class-conditional density estimation as one approach for optimal Bayesian classifier design when data availability is limited. Here we also consider the use of prior knowledge from biological pathways, along with measure transport techniques, in the computational experiments reported in Section 5. In the workflow from Figure 1, we envision prior knowledge coming from other sources as well, such as large language models [33, 34] and human experts. This is the subject of future work.
@@
The reader is referred to [14] for a thorough treatment of the topic of classification via probability density distribution functions (kernel density classification, per the terminology from [14]), among many other classification methods. We summarize next what is needed to make the discussions in Sections 4.2 and 5 (appearing later in the manuscript) self-contained. Assume we have $K$ classes to infer sample membership from. Upon computing, separately, an estimated probability density distribution function $\pi_{k}$ for each class $k$, where $k=0,\ldots,K-1$, the probability $\mathrm{Pr}(C=k\,|\,\mathcal{{Y}}=y)$ that a given sample $y$ belongs to class $k$ can be obtained from the class-conditional densities $\{\pi_{l}\}_{l=0}^{K-1}$ via
$\mathrm{Pr}(C=k\,|\,\mathcal{{Y}}=y)$ $=$ $\frac{\widehat{\pi}_{k}\,\pi_{k}(y)}{\sum_{l=0}^{K-1}\widehat{\pi}_{l}\,\pi_{l}(y)}\,,$ (1)
where $\{\widehat{\pi}_{l}\}_{l=0}^{K-1}$, $\sum_{l=0}^{K-1}\widehat{\pi}_{l}=1$, are estimates for the class prior probabilities. The latter may be determined from the proportion of training samples, for instance, or other prior knowledge.
@@
Finally, in Section 5.2 we discuss how, as part of the computational framework from Figure 1, measure transport methods have the potential to address other problems, in addition to the above-mentioned classification one. More precisely, by uncovering information hidden in the data, adaptive transport maps [2] could provide a means for generating hypotheses regarding gene relationships for the radiation biology application considered in Section 5.2. Before getting into such details, we now continue with an overview of required topics from measure transport theory.
@@ 3 Basics of Measure Transport
As mentioned in the Introduction (Section 1), optimal measure transport addresses the problem of minimizing the cost of transporting one (probability) measure to another. We now give a summary of the basic concepts, as they relate to the problem of density estimation which is of interest to us. For a thorough treatment of the subject of measure transport, the reader may refer to [44, 37, 19, 4] (theory), [24, 36, 2] (computational), and references therein.
@@
Informally, given two probability measures $\nu_{\rho}$ and $\nu_{\pi}$ defined on the algebra of Borel sets of $\mathbb{R}^{m}$, $m\geq 1$, and absolutely continuous with respect to the Lebesgue measure $\mathrm{d}{{x}}$ on $\mathbb{R}^{m}$, with densities $\rho$ and $\pi$, respectively, an invertible map $T\!:{\mathbb{R}^{m}}\rightarrow{\mathbb{R}^{m}}$ transports $\nu_{\rho}$ to $\nu_{\pi}$ if and only if
$\nu_{\pi}({B})$ $=$ $\nu_{\rho}(T^{-1}({B}))$ (2)
@@
for every Borel set $B\subset\mathbb{R}^{m}$. One says the map $T$ pushes forward $\nu_{\rho}$ to $\nu_{\pi}$, denoted $T_{\sharp}\nu_{\rho}\ =\ \nu_{\pi}$. Since we assume that both probability measures $\nu_{\rho}$ and $\nu_{\pi}$ are absolutely continuous with respect to the Lebesgue (volume) measure $\mathrm{d}{{x}}$ on $\mathbb{R}^{m}$, one may identify them with their probability densities $\rho$ and $\pi$. That is, $\nu_{\rho}(\mathrm{d}{{x}})=\rho\,\mathrm{d}{{x}}$ and $\nu_{\pi}(\mathrm{d}{{x}})=\pi\,\mathrm{d}{{x}}$. Moreover, we restrict the class of maps $T$ that we consider to the class of differentiable or even infinitely differentiable invertible maps. With these assumptions one can write the measure transport equation (2) in terms of densities. Namely, the density $\rho$ is the pullback $T^{\sharp}\pi$ of the density $\pi$,
$\rho(\cdot)\ =\ T^{\sharp}\pi(\cdot)\ =\ \pi(T(\cdot))\,|\det J_{T}(\cdot)|\,,$ (3)
where $J_{T}$ is the Jacobian matrix of $T$. Conversely, the inverse $T^{-1}$ of $T$ pushes forward $\nu_{\pi}$ to $\nu_{\rho}$, denoted $(T^{-1})_{\sharp}\nu_{\pi}=\nu_{\rho}$, and $\pi$ is the pullback $(T^{-1})^{\sharp}\rho$ of $\rho$:
$\pi(\cdot)\ =\ (T^{-1})^{\sharp}\rho(\cdot)\ =\ \rho(T^{-1}(\cdot))\,|\det J_{T^{-1}}(\cdot)|\,,$ (4)
@@
where $J_{T^{-1}}$ is the Jacobian matrix of $T^{-1}$. An illustration of these fundamental concepts appears in Figure 2.
Figure 2: The map $T:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ transports samples from $\nu_{\rho}$, a multivariate standard normal (i.e., Gaussian), to samples from $\nu_{\pi}$, a multivariate, multi-modal density distribution. The density $\rho$ is the pullback $T^{\sharp}\pi$ of the density $\pi$. Conversely, the map $T^{-1}$ transports samples from $\nu_{\pi}$ to samples from $\nu_{\rho}$. The density $\pi$ is the pullback $(T^{-1})^{\sharp}\rho\equiv T^{\sharp}\rho$ of the density $\rho$. Contour plots of the corresponding one- and two-dimensional marginal distribution densities also are shown.

### v4

@@ Abstract
One among several advantages of measure transport methods is that they allow for a unified framework for processing and analysis of data distributed according to a wide class of probability measures. Within this context, we present results from computational studies aimed at assessing the potential of measure transport techniques, specifically, the use of triangular transport maps, as part of a workflow intended to support research in the biological sciences. Scenarios characterized by the availability of limited amount of sample data, which are common in domains such as radiation biology, are of particular interest. We find that when estimating a distribution density function given limited amount of sample data, adaptive transport maps are advantageous. In particular, statistics gathered from computing series of adaptive transport maps, trained on a series of randomly chosen subsets of the set of available data samples, leads to uncovering information hidden in the data. As a result, in the radiation biology application considered here, this approach provides a tool for generating hypotheses about gene relationships and their dynamics under radiation exposure.
Keywords: density estimation, computational measure transport, triangular transport maps, gene expression data, biostatistics, statistical learning
@@ 1 Introduction
The problem of estimating a probability distribution density from samples (e.g., observations, measurements, or simulation data) is ubiquitous in data science, uncertainty quantification, clustering and classification, and probabilistic modeling and inference tasks. Moreover, it is common among various scientific and engineering fields, including biology [38, 14, 1, 41, 5, 7, 12, 40, 43]. Often, well-known parametric density functions (dependent on few parameters), such as the Gaussian or Weibull density distribution functions, are adopted. While this may simplify certain tasks (e.g., computational ones), many of these known density distribution functions are not necessarily suitable for characterizing data that exhibit complex features, such as (spatial and/or temporal) correlations and non-Gaussian characteristics.
Furthermore, in practice it is often not known beforehand which is the underlying probability distribution density that typifies a given set of samples. Having a common density estimation framework, that can seamlessly process Gaussian and non-Gaussian data alike, is thus clearly advantageous. For instance, as reported in [7], accounting for differences in the distribution densities of gene expressions can lead to improved interpretation of cancer transcriptomic data. This is just one situation (and one can envision many) where it would be beneficial to avoid the task of having to “hand-pick”, for each of a number of different data sets, the best underlying density among a list of candidate (parametric or non-parametric) distribution densities. Hence, a unified density estimation framework capable of characterizing a diverse range of properties is highly desirable. A measure transport approach [45, 37, 36] offers this possibility.
@@
Optimal measure transport, broadly defined, deals with the problem of minimizing the cost of transporting one (probability) measure to another. The subject has a long history with significant impact in many fields of mathematics, such as probability theory, optimization, and partial differential equations, among others [45, 37]. Current interest on the topic stems from its potential in numerous areas, including density estimation, data science, machine learning, Bayesian inference, sampling, and data assimilation [1, 42, 41, 24, 8, 39, 36, 32, 11, 40, 46, 17]. Among various measure transport techniques under active research, we mention the use of triangular transport maps, normalizing flows, gradient flows, diffusion maps, and invertible neural networks. Although differing in the particulars, affording a unified framework for processing and analysis of data from a range of distributions (Gaussian, non-Gaussian, uni-modal, multi-modal, etc.) is a powerful property common to all.
In this paper we consider measure transport for density estimation from data, in particular the use of triangular transport maps [19, 4, 24, 3], envisioning its potential within a computational workflow for tasks associated with biological applications. Contributions of this paper are summarized in Section 1.1 below. As described briefly next and in more detail in Sections 3 and 4, adoption of a measure transport framework for density estimation (given data samples) provides several advantages. We have already mentioned that it avoids the need to “hand-pick” among a list of candidate densities and that it can be used, seamlessly, for Gaussian and non-Gaussian data alike. In addition:
  • It yields an explicit formula for the unknown probability distribution density we wish to estimate. As will be described in Sections 3 and 4, once a suitable transport map is computed, or learned, evaluation of the unknown density at a given sample is simply done via said formula. (This will be defined precisely in Section 3, equation (7); see also equations (2)–(3) below.)
  • Generative modeling capabilities are enabled in the unknown probability space because, via the learned transport map, one can transport samples from a known probability measure to the unknown probability measure of interest.
  • If new samples from the unknown measure become available, data assimilation is facilitated without the need to learn (or re-train) from scratch the density estimation model (which, within the present context, is the transport map).
Moreover, some techniques, such as the adoption of triangular transport maps, allow one to infer information regarding the dependence structure among random variables in the unknown probability space. We make use of this property in one methodology we propose and develop, which is briefly described further below in Section 1.1 and discussed in detail in Section 5.2.2.
The mathematical foundations of measure transport, as they relate to the problem of density estimation we are concerned with, will be discussed in Sections 3 and 4. Here we introduce the setting, in a nutshell. Let $\mathcal{{Y}}=\left(\mathcal{{Y}}_{1},\mathcal{{Y}}_{2},\dots,\mathcal{{Y}}_{m}\right)\in\mathbb{R}^{m}$ be a vector of $m$ random variables jointly distributed according to an unknown probability measure $\nu_{\psi}$ defined on the $\sigma$-algebra $\mathcal{B}(\mathbb{R}^{m})$ of Borel sets of $\mathbb{R}^{m}$ and absolutely continuous with respect to the Lebesgue measure. Let $\psi$ denote the (unknown) distribution density function of $\nu_{\psi}$. It is the unknown measure $\nu_{\psi}$ and its density $\psi$ that we wish to estimate. We thus refer to $\nu_{\psi}$ and $\psi$, respectively, as the target measure and target density. Under the assumption that $\nu_{\psi}$ is absolutely continuous with respect to the Lebesgue measure, [^4] As will be discussed in Section 3, we consider only measures that are absolutely continuous with respect to the Lebesgue measure. However, the subject of measure transport is not restricted to such. It is outside the scope of this paper to treat other types of measures (such as discrete ones), but the reader may refer to [45, 37, 36] and references therein for more information. one may identify $\nu_{\psi}$ with its probability density $\psi$. In other words, we have
$\int_{B}\nu_{\psi}(\mathrm{d}{{y}})$ $=$ $\int_{B}\psi(y)\,\mathrm{d}{{y}}\,,$ (1)
for every $B\in\mathcal{B}(\mathbb{R}^{m})$. Say we are given a set $Y=\{y\in\mathbb{R}^{m}\,:\,y\sim\nu_{\psi}\}$ of samples $y\in\mathbb{R}^{m}$, which are values assumed by the random vector $\mathcal{{Y}}=\left(\mathcal{{Y}}_{j}\right)_{j=1}^{m}$. The general problem is to estimate the unknown target measure $\nu_{\psi}$, given the set of samples $Y$. By (1), to estimate the unknown target measure $\nu_{\psi}$, it thus suffices to estimate its unknown probability density function $\psi$.
Now, to estimate the unknown target probability density function $\psi$ using a measure transport approach, let $\mathcal{{X}}=\left(\mathcal{{X}}_{1},\mathcal{{X}}_{2},\dots,\mathcal{{X}}_{m}\right)\in\mathbb{R}^{m}$ be a vector of $m$ random variables jointly distributed according to a known probability measure $\nu_{\rho}$ on $\mathcal{B}(\mathbb{R}^{m})$ (assumed to be absolutely continuous with respect to the Lebesgue measure). Call $\nu_{\rho}$ the reference measure and let $\rho$ represent the (known) probability density function of $\nu_{\rho}$. As will be discussed in Sections 3 and 4, for practical purposes the reference measure $\nu_{\rho}$ should be chosen so that both evaluating its density $\rho$ and drawing samples from $\nu_{\rho}$ can be done efficiently. A common candidate for the reference measure is the $m$-variate standard normal (i.e., Gaussian) measure $\mathcal{N}(0,I)$, and that is our choice in the present work.
The central problem then becomes to compute, or learn, an invertible map $T\!:\!\mathbb{R}^{m}\!\rightarrow\!\mathbb{R}^{m}$ that “couples” the measures $\nu_{\rho}$ and $\nu_{\psi}$, where the meaning of the term “couples” will be made precise, mathematically, in Section 3 (equation (5)). Here is suffices to note that the map $T$ and its inverse $T^{-1}$ transport samples between $\nu_{\rho}$ and $\nu_{\psi}$. That is, $T$ maps any given sample $x$ from $\nu_{\rho}$ to a sample $y=T(x)$ from $\nu_{\psi}$ and, conversely, $T^{-1}$ maps any given sample $y$ from $\nu_{\psi}$ to a sample $x=T^{-1}(y)$ from $\nu_{\rho}$. Furthermore, the measure transport framework yields explicit formulas for the densities (more on this in Section 3, equations (6)–(7); also (2)–(3) below). These fundamental concepts are illustrated pictorially in Figure 1 and expanded upon in Section 3. Details regarding the mathematical representation and numerical computation of transport maps will be provided in Section 4 as well as in Appendix A.
Figure 1: The map $T\!:\!\mathbb{R}^{3}\!\rightarrow\!\mathbb{R}^{3}$ transports samples from $\nu_{\rho}$, a three-variate standard normal (i.e., Gaussian) measure $\mathcal{N}(0,I)$ with density $\rho(x)=(2\pi)^{-3/2}\,e^{{-||x||^{2}}/2}$, to samples from $\nu_{\psi}$, a three-variate measure, with multi-modal density $\psi$. The density $\rho$ is the pullback $T^{\sharp}\psi$ of the density $\psi$, and so we have $\rho(x)=T^{\sharp}\psi(x)=\psi(T(x))\,\det(J_{T}(x))$, where $J_{T}$ is the Jacobian matrix of $T$. Conversely, the map $T^{-1}$ transports samples from $\nu_{\psi}$ to samples from $\nu_{\rho}$. The density $\psi$ is the pullback $(T^{-1})^{\sharp}\rho$ of the density $\rho$, and so $\psi(y)=(T^{-1})^{\sharp}\rho(y)=\rho(T^{-1}(y))\,\det(J_{T^{-1}}(y))$. Contour plots of the corresponding one- and two-dimensional marginal distribution densities also are shown.
As will be discussed in Section 3, within the measure transport framework the unknown target density $\psi$ of interest to us is given explicitly by the formula
$\psi(y)$ $=$ $\rho(T^{-1}(y))\,\det(J_{T^{-1}}(y))\,,$ (2)
where $J_{T^{-1}}$ is the Jacobian matrix of $T^{-1}$ and $\rho$ is the chosen, known reference density function. In particular, when the reference measure $\nu_{\rho}$ is chosen to be the $m$-variate standard normal (i.e., Gaussian) measure $\mathcal{N}(0,I)$ with density $\rho(x)=(2\pi)^{-m/2}\,e^{-||x||^{2}/2}$, where $||\boldsymbol{\cdot}||$ denotes the Euclidean norm on $\mathbb{R}^{m}$, formula (2) for the target density $\psi$ becomes
[…]
Hence, once the transport map $T$ is learned, given a (new) sample $y$, the value $\psi(y)$ of the target density is thus obtained via equation (2), since all terms on the right-hand-side of (2) are known and computable. In addition, as previously noted, a sample $x$ drawn from the (known) reference measure $\nu_{\rho}$ yields a sample $y=T(x)$ from the (unknown) target measure $\nu_{\psi}$. If the reference measure $\nu_{\rho}$ is purposefully chosen so that sampling from it can be done efficiently, and since one can now transport samples from the reference to the target via the learned transport map $T$, samples from the target measure $\nu_{\psi}$ can be generated efficiently as well (as long as the transport map $T$ can be evaluated efficiently). Hence, by construction, tasks requiring probabilistic modeling, inference, and statistical analysis in the target space are enabled via a measure transport framework.
1.1 Contributions of this Paper
As previously noted, we consider the problem of measure transport for density estimation from data, in particular the use of triangular transport maps, envisioning its potential within a computational workflow for tasks associated with biological applications. The context is density estimation (or prior/posterior construction) when the number of data samples for model training is limited in amount. More precisely, we are concerned with sizes of data sets of order $O(1)$$O(10)$ samples, which are common in the radiation biology domain, yet pose challenges when faced with the problem of estimating distribution densities from data (whether or not measure transport methods are employed for the purpose). The motivation for our work will be described in Section 2, along with the proposed computational workflow for biological research. We note, however, that the methods we investigate and develop are not restricted to the particular application motivating our research.
We show that when estimating a probability density function given limited amount of sample data for model training, and without any a priori assumptions about the structure of the target probability space, adaptive transport maps [3] are advantageous. As will be explained in Section 5.2, due to the small number ($O(1)$$O(10)$) of samples at our disposal for model training, it was not sufficient to rely on learning a single (adaptive) transport map using the totality of samples available for model training. However, we introduce a method whereby statistics gathered from learning series of adaptive transport maps, trained on a series of randomly chosen subsets of the set of available data samples, offers a viable alternative in robustly revealing the structure of the target measure by distinguishing a statistically dominant triangular transport map with a specific dominant component structure. Precisely, we find that with our proposed randomized technique it is indeed possible to infer information about the dependence structure among the random variables in the unknown (multi-variate) target space. As a result, in the radiation biology application considered in Section 5.2, this approach provides a tool for generating hypotheses about gene relationships and their dynamics under radiation exposure. Details are expanded upon in Section 5.2.2 and future research directions stemming from our studies are discussed in Section 6.
The measure transport techniques we consider (and to be described in Sections 34) are at present actively researched but, to the best of our knowledge, there are few studies in the literature that have examined scenarios with the limited number of samples for model training which characterizes our motivating application. The adaptive transport maps approach proposed in [3] was shown by its authors to perform well for density estimation problems when training on data sets of various sizes (mostly $O(100)$ samples or larger). In particular, good performance was reported for data sets with $O(100)$ samples, when compared to the conditional kernel density estimation method and the $\epsilon$-neighborhood kernel density estimation method, using four data sets from the UC Irvine Machine Learning Repository [18]. It is thus one of the measure transport techniques we evaluate and build upon for our purposes.
1.2 Organization of the Paper
The rest of the paper is organized as follows. In Section 2 we introduce the application motivating our research. A discussion of the foundational mathematical and computational concepts from the subject of measure transport is provided in Sections 34, as well as in Appendix A. The mathematical theory is covered in Section 3, whereas Section 4 and Appendix A focus on numerical methods and computational aspects. Section 4.2 includes generic examples (i.e., not specific to the application motivating our work) which illustrate the measure transport framework for density estimation, seamlessly used to process both Gaussian and non-Gaussian data sets.
In Section 5 and the accompanying Appendix B we present results from our computational experiments, aimed at assessing the performance of select measure transport techniques for density estimation for a gene expression data set from the radiation biology domain. Sections 5.1 and 5.2.1 focus on phenotypic classification using measure transport densities estimated from the data set, whereas Section 5.2.2 is concerned with the problem of inferring the dependence structure between random variables (which, within the context, are associated with genes) in the unknown target probability space. (In other words, the problem of inferring the structure of the unknown target measure and its marginals.) Finally, Section 6 contains a discussion of our results and future research directions.
@@ 2 Motivating Application
The present work stems from the conception of a comprehensive computational workflow intended to support research in the biological sciences, in particular the radiation biology domain. Within this effort, our interest is in enabling a unified framework where probabilistic modeling, inference, and statistical analysis can be carried out for a wide range of data distributions. As noted in the Introduction (Section 1), measure transport offers the possibility of such unified approach. We envision measure transport techniques comprising one component within the larger scope, as depicted in Figure 2.
@@
Figure 2: Computational workflow for probabilistic modeling, inference, and statistical analysis. To account for availability of limited number of data samples, model training can be augmented with prior knowledge. Such prior knowledge may come from large language models or human experts, for instance. Measure transport techniques allow for a unified framework for density estimation and processing of data exhibiting diverse characteristics.
@@
One motivating problem within the workflow from Figure 2 is that of probability distribution density estimation for phenotypic classification problems, for example, for inferring class labels, such as radiation exposure levels or cancer types, given sample data. Normalizing flows (one kind among the various measure transport methods) have been used successfully for clustering and classification in the cancer domain [1]. Now, when the quantity of data samples available for model training is limited in amount, regardless of the density estimation technique adopted, prior knowledge (associated with the particular application) may be used in conjunction with data to estimate the class-conditional densities. In the work [5], the authors proposed incorporation of prior knowledge from biological pathways for class-conditional density estimation as one approach for optimal Bayesian classifier design when data availability is limited. Here we also consider the use of prior knowledge from biological pathways, along with measure transport techniques, in the computational experiments reported in Sections 5.1 and 5.2.1 for phenotypic classification problems, for which we employ the classification technique described shortly below (via equation (4) and the associated explanation). In the workflow from Figure 2, we envision prior knowledge coming from other sources as well, such as large language models [33, 34] and human experts. This is the subject of future work.
It is outside the scope of the present paper to cover the subject of classification itself and the various techniques available for classifying samples into different classes. The reader is referred to [14] for a thorough treatment of the topic of classification via probability density distribution functions (kernel density classification, per the terminology from [14]), among many other classification methods. Following [14], we summarize next the technique we employed for the classification results reported in Sections 4.2, 5.1, and 5.2.1. The particular technique was selected because it relies on evaluation of class-conditional densities given samples and thus fits into our proposed framework, for which explicit formulas for the class-conditional densities would be available for evaluation (refer to equation (2) and the associated discussion).
@@
Assume we have $K$ classes to infer sample membership from. Upon computing, separately (and with any suitable density estimation method), an estimated probability density distribution function $\psi_{k}$ for each class $k$, where $k=0,\ldots,K-1$, the probability $\mathrm{Pr}(C=k\,|\,\mathcal{{Y}}=y)$ that a given sample $y$ belongs to class $k$ can be obtained from the class-conditional densities $\{\psi_{l}\}_{l=0}^{K-1}$ via
$\mathrm{Pr}(C=k\,|\,\mathcal{{Y}}=y)$ $=$ $\frac{\widehat{\psi}_{k}\,\psi_{k}(y)}{\sum_{l=0}^{K-1}\widehat{\psi}_{l}\,\psi_{l}(y)}\,,$ (4)
where $\{\widehat{\psi}_{l}\}_{l=0}^{K-1}$, $\sum_{l=0}^{K-1}\widehat{\psi}_{l}=1$, are estimates for the class prior probabilities. The estimates for the class prior probabilities may be determined from the proportion of training samples, for instance, or from other prior knowledge. In Sections 4.2, 5.1, and 5.2.1, we compare our classification results with those obtained using a naive Bayes classifier, for which the class-conditional densities in (4) were estimated by fitting the training data to a Gaussian distribution, and with a support vector machine classifier, an effective classification method (not for density estimation) that constructs hyperplanes that approximate the decision boundaries separating the different underlying classes in order to perform classification of samples [14].
@@
Finally, in Section 5.2.2 we discuss how, as part of the computational framework from Figure 2, measure transport methods have the potential to address other problems, in addition to the above-mentioned classification one. More precisely, by using the randomization technique we propose in Section 5.2.2 to infer the dependence structure of the underlying unknown probability measure (i.e., dependence relations between the random variables), adaptive transport maps [3] could provide a means for generating hypotheses regarding gene relationships for the radiation biology application under consideration. Before getting into such details, we now continue with an overview of required topics from measure transport theory.
@@ 3 Basics of Measure Transport
As mentioned in the Introduction (Section 1), optimal measure transport addresses the problem of minimizing the cost of transporting one probability measure to another. [^5] Recall from Section 1 that we consider only measures that are absolutely continuous with respect to the Lebesgue measure. We now give a summary of the basic concepts, as they relate to the problem of density estimation which is of interest to us. For a thorough treatment of the subject of measure transport, the reader may refer to [45, 37, 19, 4] (theory), [24, 36, 3] (computational), and references therein.
@@
Informally, given two probability measures $\nu_{\rho}$ and $\nu_{\psi}$ defined on the $\sigma$-algebra of Borel sets of $\mathbb{R}^{m}$, $m\geq 1$, and absolutely continuous with respect to the Lebesgue measure $\mathrm{d}{{x}}$ on $\mathbb{R}^{m}$, with densities $\rho$ and $\psi$, respectively, an invertible map $T\!:{\mathbb{R}^{m}}\rightarrow{\mathbb{R}^{m}}$ transports $\nu_{\rho}$ to $\nu_{\psi}$ if and only if
$\nu_{\psi}({B})$ $=$ $\nu_{\rho}(T^{-1}({B}))$ (5)
for every Borel set $B\subset\mathbb{R}^{m}$. One says the map $T$ pushes forward $\nu_{\rho}$ to $\nu_{\psi}$, denoted $T_{\sharp}\nu_{\rho}\ =\ \nu_{\psi}$. Given any sample $x\in\mathbb{R}^{m}$ from $\nu_{\rho}$, $T$ maps $x$ to a sample $y=T(x)$ from $\nu_{\psi}$. Conversely, the inverse $T^{-1}$ of $T$ pushes forward $\nu_{\psi}$ to $\nu_{\rho}$, denoted $(T^{-1})_{\sharp}\nu_{\psi}=\nu_{\rho}$, and so $T^{-1}$ maps any given sample $y\in\mathbb{R}^{m}$ from $\nu_{\psi}$ to a sample $x=T^{-1}(y)$ from $\nu_{\rho}$.
@@
Now, since we assume that both probability measures $\nu_{\rho}$ and $\nu_{\psi}$ are absolutely continuous with respect to the Lebesgue (volume) measure $\mathrm{d}{{x}}$ on $\mathbb{R}^{m}$, one may identify them with their probability densities $\rho$ and $\psi$. That is, $\nu_{\rho}(\mathrm{d}{{x}})=\rho\,\mathrm{d}{{x}}$ and $\nu_{\psi}(\mathrm{d}{{x}})=\psi\,\mathrm{d}{{x}}$ (cf. equation (1)). Moreover, we restrict the class of maps $T$ that we consider to the class of differentiable or even infinitely differentiable invertible maps. With these assumptions one can write the measure transport equation (5) in terms of densities. Namely, the density $\rho$ is the pullback $T^{\sharp}\psi$ of the density $\psi$,
$\rho(\cdot)\ =\ T^{\sharp}\psi(\cdot)\ =\ \psi(T(\cdot))\,\det(J_{T}(\cdot))\,,$ (6)
where $J_{T}$ is the Jacobian matrix of $T$. Conversely, the density $\psi$ is the pullback $(T^{-1})^{\sharp}\rho$ of the density $\rho$:
$\psi(\cdot)\ =\ (T^{-1})^{\sharp}\rho(\cdot)\ =\ \rho(T^{-1}(\cdot))\,\det(J_{T^{-1}}(\cdot))\,,$ (7)
@@
where $J_{T^{-1}}$ is the Jacobian matrix of $T^{-1}$. An illustration of these fundamental concepts appeared previously in Section 1, Figure 1.